# LagCeresa.bib

### v1

@@ 6. Relating tropical and symplectic fluxes
Given a tropical manifold $B$, one can consider the Lagrangian torus fibration
$X(B):=T^{*}B/T_{\mathbb{Z}}^{*}B\rightarrow B,$
with the symplectic structure induced by the canonical one on $T^{*}B$. If $B$ is a torus, this fibration is trivial, and because furthermore the local system $T_{\mathbb{Z}}^{*}B$ is constant, it carries a canonical trivialisation
@@
where $F_{b}:=(T^{*}B)_{b}/(T^{*}_{\mathbb{Z}}B)_{b}$. This is given by parallel transporting elements in any fibre $F_{b^{\prime}}$ for $b^{\prime}\in B$ to the fibre $F_{b}$ by the flat connection determined by the affine structure on $B$. Because there is no ambiguity, we will usually omit the subscript $b$ and simply write $F\cong T^{n}$ for any fibre of $X(B)$.
The product decomposition above implies a Künneth decomposition of the homology of $X(B)$ into terms of the form $H_{j}(B)\otimes H_{k}(F)$. Each of these is in fact isomorphic to the tropical homology groups $H_{j}(B;\Lambda^{n-k}T_{\mathbb{Z}}B)\cong H_{j}(B;\mathbb{Z})\otimes\Lambda^{n-k}T_{\mathbb{Z}}B$ through
(6.2) $\Lambda^{n-k}T_{\mathbb{Z}}B\cong(\Lambda^{n-k}T_{\mathbb{Z}}^{*}B)^{*}\cong H_{n-k}(F;\mathbb{Z})^{*}\cong H^{n-k}(F;\mathbb{Z})\cong H_{k}(F;\mathbb{Z}),$
where the last isomorphism is Poncaré duality, and they are all canonical after a choice of orientation on the fibres. An identical reasoning holds at the level of cohomology, where we find
$H^{j}(B)\otimes H^{k}(F)\cong H^{j}(B;\Lambda^{n-k}T_{\mathbb{Z}}^{*}B)$
through the chain of isomorphisms dual to 6.2, that is
(6.3) $\Lambda^{n-k}T_{\mathbb{Z}}^{*}B\cong H_{n-k}(F;\mathbb{Z})\cong H^{k}(F;\mathbb{Z}).$
These isomorphisms of (co)homology groups are the key to relating tropical fluxes in $B$ in the sense of Section 5.3 to symplectic fluxes in $X(B)$ in the sense of Section 2. Nevertheless, they are not sufficient, and one needs to explicitly realise them at the level of (co)chains. One reason for this is to understand what differential forms on $X(B)$ the determinantal forms are mapped to. Another is to show that tautological tropical cycles $\overline{Z}$ are mapped to something resembling a PL Lagrangian lift of $Z$, as well as understand the behaviour of chains of the form $\tilde{W}$ constructed in 5.1 under these maps.
@@
It is natural to wonder how much of our construtions extend to more general settings, for instance when $B$ is singular, or the local system $T_{\mathbb{Z}}B$ has non-trivial monodromy. In the former situation, the constructions from this section do not translate, and our arguments are no longer valid globally. On the other hand, in a Lagrangian torus fibration which is non-singular, but with non-trivial monodromy for the local systems $T_{\mathbb{Z}}^{(*)}B$, we expect that our construction can be generalised by upgrading our maps to maps from (co)homology with local coefficients. However, the more complicated global topology of $X(B)$ implies the maps on homology and cohomology constructed in the following sections would no longer be isomorphisms, hence the correspondence between tropical and symplectic would no longer be one-to-one.
@@ 6.1. Map on homology
with values in cycles satisfying $[\kappa(x)]=x$. In other words, a splitting of the short exact sequence
$\text{im}(d:C_{\bullet}(F;\mathbb{Z})\rightarrow C_{\bullet}(F;\mathbb{Z}))\hookrightarrow\text{ker}(d:C_{\bullet}(F;\mathbb{Z})\rightarrow C_{\bullet}(F;\mathbb{Z}))\twoheadrightarrow H_{\bullet}(F;\mathbb{Z}),$
which exists because $F\cong T^{n}$.
By 6.2, we have isomorphisms
$C_{\bullet}(B;\Lambda^{k}T_{\mathbb{Z}}B)\cong C_{\bullet}(B;\mathbb{Z})\otimes\Lambda^{k}T_{\mathbb{Z}}B\cong C_{\bullet}(B;\mathbb{Z})\otimes H_{n-k}(F;\mathbb{Z}).$
Therefore, given a $\kappa$ as above and the trivialisation $t$ 6.1, we can define a chain map
@@
$C_{\bullet}(B;\mathbb{Z})\otimes H_{\bullet}(F;\mathbb{Z})\stackrel{{\scriptstyle(id,\kappa)}}{{\longrightarrow}}C_{\bullet}(B;\mathbb{Z})\otimes C_{\bullet}(F;\mathbb{Z})\stackrel{{\scriptstyle\text{quism}}}{{\longrightarrow}}C_{\bullet}(B\times F;\mathbb{Z})\stackrel{{\scriptstyle(t^{-1})_{*}}}{{\longrightarrow}}C_{\bullet}(X(B);\mathbb{Z}).$
The fact that this is a chain map with the differential on $C_{\bullet}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ given by restricting the coefficients for the usual singular differential is a straightforward verification, and the fact that the induced map on homology doesn’t depend on the choice of $\kappa$ follows from the condition that $[\kappa(x)]=x$. Therefore we have a well-defined map
@@
It is instructive to understand this map more explicitly. For this, we construct a choice of $\kappa$, and study the associated map $\hat{\phi}$ at the level of chain complexes. Let $\beta$ be an element in $\Lambda^{k}T_{\mathbb{Z}}B\cong H_{n-k}(F;\mathbb{Z})\cong\Lambda^{n-k}T_{\mathbb{Z}}^{*}B$. From this perspective, it can be written (although not uniquely) as
(6.5) $\beta=\sum_{i=1}^{s}e_{i,1}\wedge\dots\wedge e_{i,n-k}\in\Lambda^{n-k}T_{\mathbb{Z}}^{*}B$
for some $s\in\mathbb{N}$ and covectors $e_{i,l}\in T^{*}_{\mathbb{Z}}B$, where $l\in\{1,\dots,n-k\}$. Choosing such a decomposition amounts to assigning to each element in $H_{n-k}(F;\mathbb{Z})$ a set of $s$ finite sets of $n-k$ covectors. To these $s$ finite sets we associate $s$ parallelepipeds in $T_{\mathbb{Z}}^{*}B$, defined for $i\in\{1,\dots,s\}$ as
$E_{i}:=\{(b,t_{1}e_{i,1}+t_{2}e_{i,2}+\dots+t_{n-k}e_{i,n-k}):(t_{1},\dots,t_{n-k})\in[0,1]^{n-k}\}\subset T^{*}B.$
Passing to the quotient, this yields a cycle $E_{\beta}$ in $F$ whose homology class is $\beta$. Therefore such a choice of decomposition 6.5 yields a map $\kappa$ as in 6.4. Let $\sigma:\Delta^{j}\longrightarrow B$ be a $j$-simplex in $B$. At the chain level, $\hat{\phi}$ maps $\sigma\otimes\beta$ to $(t^{-1})_{*}(\sigma\times E_{\beta})$ in $C_{\bullet}(X(B);\mathbb{Z})$.
@@ 6.2. Map on cohomology
$\kappa^{\prime}:H^{\bullet}(F;\mathbb{Z})\longrightarrow C^{\bullet}(F;\mathbb{Z})$
with $[\kappa^{\prime}(x)]=x$. This is equivalent to assigning to each element in $\Lambda^{k}T_{\mathbb{Z}}^{*}B$ a linear combination of closed $(n-k)$-forms $dx_{i_{1}}\wedge\dots\wedge dx_{i_{n-k}}$, where $\{x_{1},\dots,x_{n}\}$ are coordinates on $B$.
@@
$C^{\bullet}(B;\mathbb{Z})\otimes H^{\bullet}(F;\mathbb{Z})\stackrel{{\scriptstyle(id,\kappa^{\prime})}}{{\longrightarrow}}C^{\bullet}(B;\mathbb{Z})\otimes C^{\bullet}(F;\mathbb{Z})\stackrel{{\scriptstyle\text{quism}}}{{\longrightarrow}}C^{\bullet}(B\times F;\mathbb{Z})\stackrel{{\scriptstyle t^{*}}}{{\longrightarrow}}C^{\bullet}(X(B);\mathbb{Z}).$
@@
Very concretely, let $\alpha$ be a $j$-form on $B$, and $\beta\in\Lambda^{k}T_{\mathbb{Z}}^{*}B$. Then using 6.3, $\beta$ can be represented by an $(n-k)$-form $\kappa^{\prime}(\beta)$ on $F$. It follows that
@@ 6.3. The pairing is preserved
We prove that the tropical fluxes we compute on the base $B$ lift to symplectic fluxes in $X(B)$. It follows from the fact that the chains of isomorphisms 6.2 and 6.3 are dual to one another that the maps $\phi$ and $\psi$ preserve the pairings in the sense that given $a$ in $H_{j}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ and $\alpha$ in $H^{j}(B;\Lambda^{k}T_{\mathbb{Z}}^{*}B)$,
@@
Let $c\in C_{j}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ and $\delta\in C^{j}(B;\Lambda^{k}T_{\mathbb{Z}}^{*}B)$. For any choices of $\kappa$ and $\kappa^{\prime}$ as in 6.1 and 6.2, the corresponding maps $\hat{\phi}$ and $\hat{\psi}$ satisfty
@@
Consider an element of $C_{j}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ of the form $\beta\otimes\sigma$, where $\sigma$ is a singular $j$-chain in $B$, and $\beta$ is viewed as an element of $H_{n-k}(F;\mathbb{Z})$, to which $\kappa$ associates a choice of representative by parallelepipeds in the fibres as in Section 6.1. Similarly, let $\alpha\otimes\eta$ be an element in $C^{j}(B;\Lambda^{k}T_{\mathbb{Z}}^{*}B)$, where $\eta$ is a $j$-form on $B$, and $\alpha$ is viewed as an element of $H^{n-k}(F;\mathbb{Z})$, to which $\kappa^{\prime}$ associates a particular choice of closed $(n-k)$-form on the fibre $F$. The tropical pairing of $\alpha\otimes\eta$ with $\beta\otimes\sigma$ is
@@
Here we have reused the notation from section 6.1, where $E_{\beta}\in C_{j+n-k}(B\times F_{b};\mathbb{Z})$ is the product of $\sigma$ with the quotient by $(T_{\mathbb{Z}}^{*}B)_{b}$ of the parallelepipeds in $T^{*}_{b}B$. It is also true that the pairing $\alpha(\beta)$ of $(n-k)$ (co)homology classes, coincides with the pairing between $\Lambda^{k}T_{\mathbb{Z}}^{*}B$ and $\Lambda^{k}T_{\mathbb{Z}}B$ through the identifications 6.2 and 6.3, which are dual through the canonical isomorphism
@@
therefore the pairing is preserved. One readily verifies that this is independent of the choice of $\kappa$ and $\kappa^{\prime}$. ∎

### v2

@@ 5. Relating tropical and symplectic fluxes
Recall from the Introduction that our proof of Theorem 1.1 involves using Zharkov’s tropical flux $\int_{\gamma_{0}}\Omega_{0}$ between $C$ and $C^{-}$ in $J(C)$ to exhibit a non-trivial symplectic flux $\int_{\hat{\phi}(\gamma_{0})}\omega^{2}$ between $L_{C}$ and $L_{C}^{-}$ in $X(J(C))$. Here $\Omega_{0}$ is the notation used by Zharkov to refer to the determinantal form we call $\Omega^{3}_{2,1}$ in Section 2.2.3, and $\gamma_{0}$ is a tropical chain in $C_{2}(J(C);T_{\mathbb{Z}}J(C))$ with $\partial\gamma_{0}=\overline{C}-\overline{C^{-}}$. The way to translate this into a symplectic flux in $X(J(C))$ is by a conormal-type construction taking $\gamma_{0}$ to $\hat{\phi}(\gamma_{0})$, and a dual construction taking $\Omega_{0}$ to $\omega^{2}$. The duality between these construction is what provides the equality
$\int_{\gamma_{0}}\Omega_{0}=\int_{\hat{\phi}(\gamma_{0})}\omega^{2}.$
In this Section we formalise this mechanism by which tropical fluxes in a tropical torus $B$ can be translated into symplectic fluxes in the corresponding Lagrangian torus fibration $X(B)$.
Because $B$ is a torus, this fibration is trivial, and because furthermore the local system $T_{\mathbb{Z}}^{*}B$ is constant, it carries a canonical trivialisation
@@
associated to an arbitrary choice of point $b\in B$, where $F_{b}:=(T^{*}B)_{b}/(T^{*}_{\mathbb{Z}}B)_{b}$. This is given by parallel transporting elements in any fibre $F_{b^{\prime}}$ for $b^{\prime}\in B$ to the fibre $F_{b}$ by the flat connection determined by the affine structure on $B$. In fact, recall that the local systems $T_{\mathbb{Z}}^{(*)}B$ are constant, because the linear part of the monodromy of the sheaf of affine functions $\textit{Aff}_{B}\subset C^{\infty}(B)$ vanishes (i.e., the slope of these functions is preserved by the monodromy). This implies that sections of this local system are all global, and can be identified with points in $(T^{*}_{\mathbb{Z}}B)_{b}$. With this in mind, we will repeatedly use the isomorphisms
$H^{j}(B;\Lambda^{k}T^{*}_{\mathbb{Z}}B)\cong H^{j}(B;\mathbb{Z})\otimes\Lambda^{k}(T^{*}_{\mathbb{Z}}B)_{b}$ and $H_{j}(B;\Lambda^{k}T_{\mathbb{Z}}B)\cong H_{j}(B;\mathbb{Z})\otimes\Lambda^{k}(T_{\mathbb{Z}}B)_{b}.$
The product decomposition above implies a Künneth decomposition of the homology of $X(B)$ into terms of the form $H_{j}(B)\otimes H_{k}(F_{b})$. Each of these is in fact isomorphic to the tropical homology groups $H_{j}(B;\Lambda^{n-k}T_{\mathbb{Z}}B)\cong H_{j}(B;\mathbb{Z})\otimes\Lambda^{n-k}(T_{\mathbb{Z}}B)_{b}$ through
(5.2) $(\Lambda^{n-k}T_{\mathbb{Z}}B)_{b}\cong(\Lambda^{n-k}T_{\mathbb{Z}}^{*}B)_{b}^{*}\cong H_{n-k}(F_{b};\mathbb{Z})^{*}\cong H^{n-k}(F_{b};\mathbb{Z})\cong H_{k}(F_{b};\mathbb{Z}),$
where the last isomorphism is Poincaré duality, and they are all canonical after a choice of orientation on the fibres. An identical reasoning holds at the level of cohomology, where we find
$H^{j}(B)\otimes H^{k}(F_{b})\cong H^{j}(B;\Lambda^{n-k}T_{\mathbb{Z}}^{*}B)$
through the chain of isomorphisms dual to that of Equation (5.2), that is
(5.3) $\Lambda^{n-k}(T_{\mathbb{Z}}^{*}B)_{b}\cong H_{n-k}(F_{b};\mathbb{Z})\cong H^{k}(F_{b};\mathbb{Z}).$
These isomorphisms of (co)homology groups are the key to relating tropical fluxes in $B$ in the sense of Section 2.2.4 to symplectic fluxes in $X(B)$ in the sense of Section 3. Nevertheless, they are not sufficient, and one needs to explicitly realise them at the level of (co)chains. One reason for this is to understand what differential forms on $X(B)$ the determinantal forms are mapped to. Another is to show that tautological tropical cycles $\overline{Z}$ are mapped to something resembling a PL Lagrangian lift of $Z$, as well as understand the behaviour of chains of the form $\tilde{W}$ constructed in Section 2.2.2 under these maps.
@@
One can wonder how much of our construtions extend to more general settings, for instance when $B$ is singular, or the local system $T_{\mathbb{Z}}B$ has non-trivial monodromy. In the former situation, the constructions from this section do not translate, and our arguments are no longer valid globally. On the other hand, in a Lagrangian torus fibration which is non-singular, but with non-trivial monodromy for the local systems $T_{\mathbb{Z}}^{(*)}B$, we expect that our construction can be generalised by upgrading our maps to maps from (co)homology with local coefficients. However, the more complicated global topology of $X(B)$ implies the maps on homology and cohomology constructed in the following sections would no longer be isomorphisms, hence the correspondence between tropical and symplectic would no longer be one-to-one.
@@ 5.1. Map on homology
with values in cycles satisfying $\kappa(x)]=x$. In other words, a splitting of the short exact sequence
$\text{im}(d:C_{\bullet}(F_{b};\mathbb{Z})\rightarrow C_{\bullet}(F_{b};\mathbb{Z}))\hookrightarrow\text{ker}(d:C_{\bullet}(F_{b};\mathbb{Z})\rightarrow C_{\bullet}(F_{b};\mathbb{Z}))\twoheadrightarrow H_{\bullet}(F_{b};\mathbb{Z}),$
which exists because $F_{b}\cong T^{n}$. Here and throughout, the point $b$ is an arbitrary point chosen in $B$.
By Equation (5.2), we have isomorphisms
$C_{\bullet}(B;\Lambda^{k}T_{\mathbb{Z}}B)\cong C_{\bullet}(B;\mathbb{Z})\otimes\Lambda^{k}(T_{\mathbb{Z}}B)_{b}\cong C_{\bullet}(B;\mathbb{Z})\otimes H_{n-k}(F_{b};\mathbb{Z}).$
Therefore, given a $\kappa$ as above and the trivialisation $t$ from Equation (5.1), we can define a chain map
@@
$C_{\bullet}(B;\mathbb{Z})\otimes H_{\bullet}(F_{b};\mathbb{Z})\stackrel{{\scriptstyle(id,\kappa)}}{{\longrightarrow}}C_{\bullet}(B;\mathbb{Z})\otimes C_{\bullet}(F_{b};\mathbb{Z})\longrightarrow C_{\bullet}(B\times F_{b};\mathbb{Z})\stackrel{{\scriptstyle(t^{-1})_{*}}}{{\longrightarrow}}C_{\bullet}(X(B);\mathbb{Z}),$
where the middle arrow is the Eilenberg-Zilber map, which is a quasi-isomorphism.
The fact that this is a chain map with the differential on $C_{\bullet}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ given by restricting the coefficients for the usual singular differential is a straightforward verification, and the fact that the induced map on homology doesn’t depend on the choice of $\kappa$ follows from the condition that $\kappa(x)]=x$. Therefore we have a well-defined map
@@
It is instructive to understand this map more explicitly. For this, we construct a choice of $\kappa$, and study the associated map $\hat{\phi}$ at the level of chain complexes. Let $\beta$ be an element in $\Lambda^{k}(T_{\mathbb{Z}}B)_{b}\cong H_{n-k}(F_{b};\mathbb{Z})\cong\Lambda^{n-k}(T_{\mathbb{Z}}^{*}B)_{b}$. From this perspective, it can be written (although not uniquely) as
(5.5) $\beta=\sum_{i=1}^{s}e_{i,1}\wedge\dots\wedge e_{i,n-k}\in\Lambda^{n-k}(T_{\mathbb{Z}}^{*}B)_{b}$
for some $s\in\mathbb{N}$ and covectors $e_{i,l}\in(T^{*}_{\mathbb{Z}}B)_{b}$, where $l\in\{1,\dots,n-k\}$. Choosing such a decomposition amounts to assigning to each element in $H_{n-k}(F_{b};\mathbb{Z})$ a set of $s$ finite sets of nkn-k covectors. To these $s$ finite sets, we associate $s$ parallelepipeds in $T^{*}B$, defined for $i\in\{1,\dots,s\}$ as
$E_{i}:=\{(b,t_{1}e_{i,1}+t_{2}e_{i,2}+\dots+t_{n-k}e_{i,n-k}):(t_{1},\dots,t_{n-k})\in[0,1]^{n-k}\}\subset(T^{*}B)_{b}.$
Passing to the quotient, this yields a cycle $E_{\beta}$ in $F_{b}$ whose homology class is $\beta$. Therefore such a choice of decomposition from Equation (5.5) yields a map $\kappa$ as in Equation (5.4). Let $\sigma:\Delta^{j}\longrightarrow B$ be a $j$-simplex in $B$. At the chain level, $\hat{\phi}$ maps $\sigma\otimes\beta$ to $(t^{-1})_{*}(\sigma\times E_{\beta})$ in $C_{\bullet}(X(B);\mathbb{Z})$.
Example 5.2.
[…]
$\kappa:H_{\bullet}(F_{b};\mathbb{Z})$ $\longrightarrow C_{\bullet}(F_{b};\mathbb{Z})$
$[\gamma_{1}]\hskip 11.38109pt$ $\hskip 2.84526pt\mapsto\hskip 17.07164pt\gamma_{1}$
$[\gamma_{2}]\hskip 11.38109pt$ $\hskip 2.84526pt\mapsto\hskip 17.07164pt\gamma_{2}.$
Now take
$x_{0}$ $:=\sigma\in C_{1}(B;\Lambda^{0}T_{\mathbb{Z}}B)\cong C_{1}(B;\mathbb{Z})$
$x_{1}$ $:=v_{1}\otimes\sigma\in C_{1}(B;\Lambda^{1}T_{\mathbb{Z}}B)$
$x_{3}$ $:=v_{1}\wedge v_{2}\otimes\sigma\in C_{1}(B;\Lambda^{2}T_{\mathbb{Z}}B).$
We compute the images $\hat{\phi}(x_{0})$, $\hat{\phi}(x_{1})$, and $\hat{\phi}(x_{2})$. Through the isomorphism
$\Lambda^{*}T_{\mathbb{Z}}B\cong\Lambda^{n-*}T_{\mathbb{Z}}^{*}B\cong H_{n-*}(F_{b};\mathbb{Z}),$
$1\in\Lambda^{0}T_{\mathbb{Z}}B$ is mapped to $\gamma_{1}]\wedge[\gamma_{2}]\in H_{2}(F_{b};\mathbb{Z})$, $v_{1}\in\Lambda^{1}T_{\mathbb{Z}}B$ is mapped to $\gamma_{2}]\in H_{1}(B;\mathbb{Z})$, and $v_{1}\wedge v_{2}\in\Lambda^{2}T_{\mathbb{Z}}B$ is mapped to $1\in H_{0}(F_{b};\mathbb{Z})$. Therefore, with our choice of $\kappa$, we have
$\hat{\phi}(x_{0})$ $:=(t^{-1})_{*}\left(\sigma\times(\gamma_{1}\wedge\gamma_{2})\right)\in C_{3}(X(B);\mathbb{Z})$
$\hat{\phi}(x_{1})$ $:=(t^{-1})_{*}\left(\sigma\times\gamma_{2}\right)\in C_{2}(X(B);\mathbb{Z})$
$\hat{\phi}(x_{3})$ $:=(t^{-1})_{*}\left(\sigma\right)\in C_{1}(X(B);\mathbb{Z}).$
@@ 5.2. Map on cohomology
$\kappa^{\prime}:H^{\bullet}(F_{b};\mathbb{Z})\longrightarrow C^{\bullet}(F_{b};\mathbb{Z})$
with $\kappa^{\prime}(x)]=x$. This is equivalent to assigning to each element in $\Lambda^{k}(T_{\mathbb{Z}}^{*}B)_{b}$ a linear combination of closed (nk)(n-k)-forms $dx_{i_{1}}\wedge\dots\wedge dx_{i_{n-k}}$, where $\{x_{1},\dots,x_{n}\}$ are coordinates on $B$.
@@
$C^{\bullet}(B;\mathbb{Z})\otimes H^{\bullet}(F_{b};\mathbb{Z})\stackrel{{\scriptstyle(id,\kappa^{\prime})}}{{\longrightarrow}}C^{\bullet}(B;\mathbb{Z})\otimes C^{\bullet}(F_{b};\mathbb{Z})\stackrel{{\scriptstyle\text{quism}}}{{\longrightarrow}}C^{\bullet}(B\times F_{b};\mathbb{Z})\stackrel{{\scriptstyle t^{*}}}{{\longrightarrow}}C^{\bullet}(X(B);\mathbb{Z}).$
@@
Very concretely, let $\alpha$ be a $j$-form on $B$, and $\beta\in\Lambda^{k}(T_{\mathbb{Z}}^{*}B)_{b}$. Then using Equation (5.3), $\beta$ can be represented by an (nk)(n-k)-form $\kappa^{\prime}(\beta)$ on $F_{b}$. It follows that
@@ 5.3. The pairing is preserved
We prove that the tropical fluxes we compute on the base $B$ lift to symplectic fluxes in $X(B)$. It follows from the fact that the chains of isomorphisms in Equations (5.2) and (5.3) are dual to one another that the maps $\phi$ and $\psi$ preserve the pairings in the sense that given $a$ in $H_{j}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ and $\alpha$ in $H^{j}(B;\Lambda^{k}T_{\mathbb{Z}}^{*}B)$,
@@
Let $c\in C_{j}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ and $\delta\in C^{j}(B;\Lambda^{k}T_{\mathbb{Z}}^{*}B)$. For any choices of $\kappa$ and $\kappa^{\prime}$ as in Equations (5.1) and (5.2), the corresponding maps $\hat{\phi}$ and $\hat{\psi}$ satisfty
@@
Consider an element of $C_{j}(B;\Lambda^{k}T_{\mathbb{Z}}B)$ of the form $\beta\otimes\sigma$, where $\sigma$ is a singular $j$-chain in $B$, and $\beta$ is viewed as an element of $H_{n-k}(F_{b};\mathbb{Z})$, to which $\kappa$ associates a choice of representative by parallelepipeds in the fibre direction as in Section 5.1. Similarly, let $\alpha\otimes\eta$ be an element in $C^{j}(B;\Lambda^{k}T_{\mathbb{Z}}^{*}B)$, where $\eta$ is a $j$-form on $B$, and $\alpha$ is viewed as an element of $H^{n-k}(F_{b};\mathbb{Z})$, to which $\kappa^{\prime}$ associates a particular choice of closed (nk)(n-k)-form on the fibre $F_{b}$. The tropical pairing of $\alpha\otimes\eta$ with $\beta\otimes\sigma$ is
@@
Here we have reused the notation from Section 5.1, where $E_{\beta}\in C_{j+n-k}(B\times F_{b};\mathbb{Z})$ is the product of $\sigma$ with the quotient by $(T_{\mathbb{Z}}^{*}B)_{b}$ of the parallelepipeds in $(T^{*}B)_{b}$. It is also true that the pairing $\alpha(\beta)$ of (nk)(n-k) (co)homology classes, coincides with the pairing between $\Lambda^{k}T_{\mathbb{Z}}^{*}B$ and $\Lambda^{k}T_{\mathbb{Z}}B$ through the identifications from Equations (5.2) and (5.3). These are dual through the canonical isomorphism
@@
therefore the pairing is preserved. One readily verifies this is independent of the choice of $\kappa$ and $\kappa^{\prime}$. ∎